Instrument Log 

MPL HR          C1                  Lamont, OK Optical Trailer

Date

Operator

Comments

8/9/06

CAM

BCR-01258 requests that MPLHR be removed from service and the new polarized MPL SN106 be installed In its place.  Powered MPLHR off at 1414 GMT.  Removed *MPLHR and all of its components.  Placed the polarized MPL in its place and RTS at 1614 GMT.  Will store the old laptop here on site and ship the old MPL and its components to Albert Mendoza to be upgraded to a polarized MPL.  Components shipped include: MPL SN04 (WD24164), Spectra Physics Laser Power Supply SN U3680 (WD27512), MPL Scaler and Control SN 009 (WD27543), Coherent Laser Power Supply SN 20126 (PT07704), Spectra Physics Remote Keypad SN 3508.  Old IBM laptop SN is 78-PMZDG (PT05821).  *Old MPLHR – 04  *New MPL-POL - 106

8/8/06

CM/MK

Returned to the site to complete the validation process for MPL 105.  Stopped normal data operations at 0247 GMT.  Removed the side port cover and tilted the MPL horizontal to collect overlap data.  Collected data until 0700 GMT.  Saved all relevant data files to the hard drive and also sent to mentor.  Restarted MPL 105 in normal data collection mode and RTS at 0711 GMT.

8/7/06

CM/MK

Stopped MPL 106 at 1400 GMT and slid out of the way.  Installed MPL 105 in its place to begin validation process.  Measured pulse width of 17ns at current setting of 1A and 13ns at 1.25A.  Took after pulse, dark count and background measurements.  Checked energy monitor calibration.  Powered MPL 105 on in normal data collection mode at 1954 GMT.

8/7/06

PAD

Last SDS collection was on 8/5/06 at 0820 GMT.  Found data files on C drive.  Rebooted computer at 1430 GMT.  RTS at 1445 GMT.

8/3/06

PAD

Last SDS collection was at 0520 GMT.  Found data files on C drive.  Rebooted computer at 1415 GMT.  RTS at 1430 GMT.

8/1/06

PAD

Last SDS collection was at 1020 GMT.  Found data files on C drive.  Rebooted MPL computer at 1418 GMT.  RTS at 1433 GMT.

7/31/06

PAD

Last successful collection was at 1420 GMT.  Found data files on C drive.  Rebooted MPL computer at 1831 GMT.  RTS at 1846 GMT.

7/24/06

PAD

Last SDS collection was on 7/22/06 at 0720 GMT.  Found data files on C drive.  Rebooted computer at 1431 GMT.  RTS at 1446 GMT.

7/18/06

PAD

Last SDS collection was on 7/17/06 at 1820 GMT.  Found data files on C drive.  Rebooted MPL computer at 1503 GMT.  RTS at 1518 GMT.

7/10/06

PAD

Last SDS collection was on 7/7/06 at 1620 GMT.  Found data files on C drive.  Rebooted computer at 1536 GMT.  RTS at 1551 GMT.

7/5/06

PAD

SDS is not collecting data files from MPL computer.  Found data files on C drive from 06070411 to current.  Rebooted MPL computer at 1416 GMT.

6/30/06

PAD

Last SDS collection was at 0120 GMT.  Found data files on C drive.  Rebooted computer at 1440 GMT.  RTS at 1455 GMT.

6/15/06

PAD

Last SDS collection was at 0620 GMT.  Found data files on C drive.  Rebooted MPL computer at 1332 GMT.  RTS at 1347 GMT.

6/15/06

CAM

Stopped MPL 106  program at 0128 GMT to collect horizontal data for validation.  Pointed MPL horizontal and collected data for several hours.  RTS in normal operational mode at 0614 GMT.  Note – MPL computer clock was set to the wrong time zone around 1530 GMT on 6/14, resulting in the clock being one hour fast for all validation data.  Set the clock to the correct time  zone at 0613 GMT.

6/14/06

CAM

Stopped MPL 106 program at 1615 GMT to begin validation procedures.  Checked pulse width and power of laser.  Performed energy monitor calibration checks.  Collected and saved dark count, background and after pulse data.  Restarted in normal operation mode (30m/10s/alt-pol) and RTS at 2051 GMT.

6/14/06

PAD

Last SDS collection was at 0120 GMT.  Found data files on C drive.  Rebooted computer at 1432 GMT.  RTS at 1447 GMT.

6/13/06

PAD

Last SDS collection was on 6/12/06 at 2120 GMT.  Found data files on C drive.  Rebooted MPL computer at 1503 GMT.  RTS at 1518 GMT.

6/13/06

CAM

The 2nd new MPL (SN106) has been received and needs to be installed for validation.  Instrument mentor requested the computer and data system of MPL 104 be returned to the Vendor for testing/diagnostics.  Powered system 104 down at 1740 GMT.  Installed system 106 and powered on at 1910 GMT.

6/12/06

PAD

Last SDS collection was on 6/10/06 at 0320 GMT.  Found data files on C drive.  Rebooted MPL computer at 1723 GMT.  RTS at 1738 GMT.

6/9/06

PAD

Last SDS collection was on 6/8/06 at 1320 GMT.  Found data files on C drive.  Rebooted computer at 2040 GMT.  RTS at 2055 GMT.

6/7/06

PAD

Last SDS collection was on 6/6/06 at 1620 GMT.  Rebooted MPL computer at 1422 GMT.  RTS at 1437 GMT.

6/7/06

CAM

Instrument mentor requested the MPLHR diode current be increased by 0.05 A.   The energy has been decreasing due to the aging diode.  Raised the diode current setting from 1330 mA to 1380 mA at 1900 GMT.  Noted an increase in the energy reading from around 5.6 uJ to 6.5 uJ.

6/2/06

MK

While cleaning the inner side of the new MPL532 window the RTV seal broke. Removed all old RTV from window and frame. Covered MPL532, cleaned window and reset window in new RTV. Removed MPL cover at 1521 GMT.

6/1/06

PAD

Last SDS collection was at 0020 GMT.  Rebooted computer at 1435 GMT.  RTS at 1450 GMT.

5/31/06

PAD

Last SDS collection was at 0320 GMT.  Rebooted computer at 1414 GMT.  RTS at 1429 GMT.

5/30/06

PAD

Last SDS collection was on 5/26/06 at 1320 GMT.  Found data files on C drive of computer.  Rebooted MPL computer at 1422 GMT.  RTS at 1437 GMT.

5/24/06

PAD

SDS Computer has not collected data since 0920 GMT.  Rebooted MPL computer at 1831 GMT.  RTS at 1846 GMT.

5/22/06

PAD

SDS has not collected data from MPL computer since 5/19/06 at 0520 GMT.  Rebooted computer at 1523 GMT.  RTS at 1538 GMT.

5/19/06

CAM

Instrument mentor notified Site Ops that the new MPL computer had locked up while they were doing remote system checks.  Went to the GIF trailer and found that Windows had shut down due to a problem caused by nidaq32k.sys – page_fault_in_nonpaged_area.  Rebooted the computer at 1708 GMT.  Restarted the Sigma program and RTS at 1709 GMT.

5/18/06

PAD

SDS computer has not collected data from MPL since 5/16/06.  Found data in C drive on computer.  Rebooted MPL computer at 1941 GMT

5/15/06

PAD

SDS computer has not collected data from MPL computer since 0120 GMT.  Rebooted computer at 1446 GMT.  RTS at 1501 GMT.

5/10/06

CAM

Morning status check found the new MPL SN104 in an error state.  'A/D Start Collection Error' was scrolling on the software display.  Tried stopping and restarting the software, but it then came up with MCS Card Not Found.  Powered everything down and restarted status ok at 1512 GMT.  However, checked status remotely a few minutes later and found it was scrolling 'A/D Start Collection Error' again.  This time, restarting just the software was successful.  RTS at 1545 GMT.  Using the following settings: Use Polarization: alt-pol.txt, 30m resolution, 10s averaging time.  Notified mentor. 

5/8/06

PAD

SDS computer is not collecting data from MPL computer.  Found data on C drive.  Rebooted computer at 1448 GMT.  RTS at 1503 GMT.

5/5/06

PAD

SDS computer is not collecting data from MPL computer.  Rebooted MPL computer at 1441 GMT.  RTS at 1456 GMT.

5/3/06 – 5/5/06

Mentor

Mentor on site to install the new MPL SN104.  MPL is currently operating side by side with the old MPLHR.  Operational as of 5/5.

5/2/06

PAD

SDS computer is not collecting data from MPL computer.  Rebooted MPL computer at 1412 GMT.  RTS at 1427 GMT.

4/28/06

PAD

SDS computer is not collecting data from MPL computer.  Rebooted MPL computer at 1856 GMT.  RTS at 1911 GMT.

4/21/06

PAD

Data is not being transferred to SDS computer.  Rebooted computer at 1412 GMT.  RTS at 1427 GMT.

4/17/06

PAD

Dust has accumulated on collector plate and will not blow off with canned air.  Corrector plate needs to be cleaned.  At 1904 GMT, stopped program and turned off laser.  Removed collector plate and cleaned with distilled water and isopropyl alcohol.  Installed collector plate and restarted laser and MPL program.  RTS at 1954 GMT.

4/12/06

PAD

Computer is not communicating with SDS.  Rebooted computer at 1420 GMT.  RTS at 1435 GMT.

4/10/06

PAD

Computer has stopped communicating with SDS.  Rebooted computer at 1438 GMT.  RTS at 1453 GMT.

4/6/06

PAD

The MPL is not transferring data at this time.  Restarted MPL laptop computer at 1453 GMT, all data communications were restored.

4/5/06

MK

The MPL is not transferring data at this time. Restarted MPL laptop computer at 1345 GMT, all data communications were restored.

4/3/06

PAD

Last SDS collection was on 4/2/06 at 0220 GMT.  Found data files in C: drive.  Rebooted computer at 1846 GMT.  RTS at 1901 GMT.

3/24/06

MK

SDS is not collecting data from MPLHR laptop.  Rebooted computer at 1502 GMT.  RTS at 1517 GMT.

3/23/06

PAD

Last successful SDS collection was on 3/22/06 at 1720 GMT.  Found data files on C drive.  Rebooted computer at 1503 GMT.  RTS at 1518 GMT.

3/22/06

PAD

Last successful SDS collection was on 3/21/06 at 1720 GMT.  Found data files on C drive.  Rebooted computer at 1524 GMT.  RTS at 1539 GMT.

3/8/06

PAD

Last successful collection was on 3/7/06 at 2120 GMT.  Found data files on C drive from 06030720 to current.  Rebooted computer at 1541 GMT.     

3/7/06

PAD

Last SDS collection was on 3/7/06 at 0520 GMT.  Found data files on C: drive from 2006030704 to current.  Rebooted computer at 1540 GMT.  RTS at 1555 GMT.

3/6/06

PAD

Last SDS collection was on 3/5/06 at 1220 GMT.  Found data files on C: drive from 2006030511 to current.  Rebooted computer at 1505 GMT.  RTS at 1520 GMT.

2/27/06

MK

MPL computer not transferring data. Rebooted MPL computer at 1635 GMT.

2/24/06

PAD

Last successful collection was on 2/23/06 at 1620 GMT.  Found data files on C drive from 06022315 to current.  Rebooted computer at 1515 GMT.  RTS at 1530 GMT.

2/22/06

MK

MPL computer is not communicating at this time. Rebooted MPL computer at 1614 GMT.

2/16/06

PAD

Last successful SDS collection was at 1020 GMT.  Found data files from 06021609 to current on C drive.  Rebooted computer at 1648 GMT.  RTS at 1703 GMT.

2/13/06

PAD

Last successful collection was on 2/11/06 at 0020 GMT.  Found data files on C drive from 06021023 to current.  Rebooted computer at 1525 GMT.  RTS at 1540 GMT.

1/31/06

PAD

Last successful collection was at 0720 GMT.  Found data files from 06013106 to current on C drive.  Rebooted computer at 1529 GMT.  RTS at 1539 GMT.

1/27/06

PAD

Last successful collection was at 0620 GMT.  Found data files from 06012705 to current on C drive.  Rebooted computer at 1439 GMT.  RTS at 1454 GMT.

1/23/06

PAD

Last successful collection was on 1/22/06 at 0620 GMT.  Found data files on C drive from 06012205 to current.  Rebooted computer at 1531 GMT.  RTS at 1541 GMT.

1/11/06

PAD

Last successful collection was on 1/10/06 at 1720 GMT.  Found data files from 06011016 to current on C: drive.  Rebooted computer at 1913 GMT.  RTS at 1928 GMT.

1/10/06

PAD

Last successful collection was at 0220 GMT.  Found data files from 06011001 to current on C: drive.  Rebooted computer at 1630 GMT.  RTS at 1645 GMT.

1/9/06

PAD

Last successful collection was on 1/7/06 at 0420 GMT.  Found data files from 06010703 to current on C: drive.  Rebooted computer at 1527 GMT.

1/6/06

PAD

Last successful collection was at 1220 GMT.  Found data files from 06010611 to current on C: drive.  Rebooted computer at 1538 GMT.  RTS at 1553 GMT.

1/3/06

PAD

Last successful collection was on 12/30/05 at 2020 GMT.  Found data files 05123019 to current on C: drive.  Rebooted computer at 1540 GMT.  RTS at 1555 GMT.

12/27/05

PAD

Last successful SDS collection was on 12/24/05 at 1720 GMT.  Found data files on C: drive from 05122416 to current.  Rebooted computer at 1529 GMT.  RTS at 1544 GMT.

12/19/05

PAD

Last successful SDS collection was on 12/18/05 at 1320 GMT.  Found data files on C: drive from 05121812 to current.  Rebooted computer at 1709 GMT.  RTS at 1724 GMT.

12/16/05

PAD

Last successful SDS collection was on 12/15/05 at 2020 GMT.  Found data files on C: drive from 05121519 to current.  Rebooted computer at 1553 GMT.  RTS at 1608 GMT.

12/13/05

PAD

Last successful SDS collection was on 12/13/05 at 0520 GMT.  Found files from 05121304 to current on C drive.  Rebooted computer at 1522 GMT.

12/12/05

PAD

Last successful SDS collection was on 12/11/05 at 1520 GMT.  Found files from 05121114 to current on C drive.  Rebooted computer at 1559 GMT.

12/8/05

PAD

Last successful SDS collection was on 12/7/05 at 1720 GMT.  Found data files from 05120716 to current on C drive.  Rebooted computer at 1517 GMT.  RTS at 1532 GMT.

12/7/05

PAD

Last successful SDS collection was on 12/6/05 at 0620 GMT.  Found data files from 05120605 to current on C drive.  Rebooted computer at 1554 GMT.  RTS at 1611 GMT.

11/28/05

PAD

SDS has not collected since 11/24/05 at 0100 GMT.  Found data files from 05112401 to current in C drive.  Rebooted computer at 1542 GMT.  SDS successfully collected data files.  RTS at 1557 GMT.

11/21/05

PAD

Last SDS collection was on 11/19/05 at 2320 GMT.  Found data files on computer.  Rebooted computer at 1944.  RTS at 1959 GMT.

11/17/05

PAD

Last SDS collection was on 11/16/05 at 1520 GMT.  Found data files on MPL computer from 05111614 up to current working file.  Rebooted computer at 1525 GMT.  RTS at 1534 GMT.

11/14/05

PAD

SDS last successful collection was on 11/12/05 at 1300 GMT.  Found data files on computer from 05111213 to current.  Rebooted computer at 1640 GMT.  RTS at 1655 GMT.

11/8/05

PAD

SDS last successful collection was at 1020 GMT.  Found files from 0900 to current time in C drive.  Rebooted computer at 1943 GMT.  RTS at 1958 GMT.               

10/26/05

TMG

Rebooted the MPLHR computer @ 2102 GMT due to ftp problems.  RTS at 2105 GMT.

10/24/05

TMG

The MPLHR laptop needed to be rebooted due to Serv-U not working properly.  Remotely rebooted laptop at 1626 GMT.  The laptop hung up while shutting down.  Went to the GIF trailer and power cycled laptop.  RTS at 1655 GMT.

10/17/05

PAD

Last SDS collection was on 10/15/05 at 2120 GMT.  Found data files in C drive.  Rebooted computer at 1425 GMT.  RTS at 1440 GMT.

10/14/05

PAD

SDS last successful collection was on 10/12/05 at 1320 GMT.  Found data files are on C drive.  Rebooted computer at 1600 GMT.  RTS at 1615 GMT.

10/11/05

PAD

Last SDS collection was on 10/9/05 at 1820 GMT.  Found data files 05100917 up to 05101115 in C drive.  Rebooted computer at 1529 GMT.  SDS collected files.  RTS at 1544 GMT.

10/6/05

PAD

Last successful collection was on 10/5/05 at 1320 GMT.  Found data files in C drive from 1200 GMT on 10/5/05 up to current time.  Stopped program and rebooted computer.  RTS at 1836 GMT.

9/27/05

PAD

Last successful collection was on 9/26/05 at 0420 GMT.  Computer is running and uncollected files are in drive C.  Rebooted computer at 1430 GMT.  SDS collected data files.  RTS at 1445 GMT.

9/8/05

PAD

Last successful collection was at 0020 GMT.  System continues to run and create files.  SDS system is communicating with MPL computer but it cannot find data files.  Computer department was troubleshooting problem and had tried to remotely reboot computer.  At 1405 GMT, found computer stuck trying to reboot. Power down computer and restarted.  Program restarted correctly.  Mentor has been notified of problem and is troubleshooting at this time.  RTS at 1420 GMT.

9/6/05

PAD

Error displayed:  About Time  This program has performed and illegal operation and will be shut down.  If the problem persists, contact the program vendor.  About time caused an invalid page fault in module ABOUTTIME.EXE at 017f: 0042aeeb.  Cleared error and restarted about time program.  Computer clock was slow and adjusted by 30534ms.  RTS at 1536 GMT.

9/1/05

PAD

Last successful collection was on 8/30/05 at 2120 GMT.  Computer and program appears to be operating correctly.  SDS shows no files to get with extensions .*.  Computer can be remotely accessed and SDS system is communicating with computer.  Program is saving data files and after computer is rebooted, SDS will collect those files.  Rebooted computer at 1411 GMT.  SDS successfully collected data files at 1420 GMT.  Troubleshooting with mentor and computer department at this time.  RTS at 1426 GMT. 

8/30/05

MR

The cover needs to be placed on the instrument so that a cut out can be installed on the bump out.  Installed the cover at 1635 GMT.  Removed the cover at 1738 GMT.

8/25/05

CAM

The SDS collection logs had been indicating problems with the MPL time server.  The computer department made changes to the MPL computer's time server network settings.  The computer needed to be rebooted.  Remotely rebooted the computer using Radmin at 1510 GMT.  The computer hung up while shutting down windows and did not restart.  Went to the GIF trailer and manually power cycled the computer.  The program restarted and RTS status O.K. at 1525 GMT.  No time server errors appeared in the following SDS collection log.

8/25/05

PAD

Last successful collection was at 0820 GMT.  At 1415 GMT, program appears to be running correctly.  Rebooted computer at 1421 GMT.  Notified instrument mentor.  Troubleshooting at this time.  RTS at 1430 GMT.

8/24/05

PAD

Last successful collection was on 8/22/05 at 2123 GMT.  SDS shows:  No files to get with extensions: .*.  Rebooted computer at 1358 GMT.  SDS collections returned to normal.  RTS at 1408 GMT.

8/22/05

PAD

Last successful collection was on 8/19/05 at 1620 GMT.  Experiencing SDS problems at CF so DSView page can not be used to check instrument collection status.  Instrument daily PM was done at 1531 GMT and program appears to be running correctly.  At 2044 GMT, stopped program and rebooted computer to reestablish communication with SDS.  Program restarted correctly.  RTS at 2059 GMT.

8/11/05

PAD

Last successful collection was at 1800 GMT.   Computer is running and program operating correctly.  Ron Reed informed me SDS is communicating with computer but there is no files to collect.  Stopped program and there was an Explore-MPL box up behind program on monitor.  Rebooted computer.  RTS at 1935 GMT.

7/25/05

PAD

Last successful collection was on 7/23/05 at 0220 GMT. Computer clock is 1 min. 40 sec. slow.  MPL program is running correctly.  Stopped MPL program and rebooted computer to reestablish communication.  After reboot, about time program could not find server.  Notified windows administrator.  Was advised that a bad fiber wire was found to server.  Computer was rebooted and time set at 1854 GMT.

7/14/05

PAD

Computer clock is 5 seconds slow.  Started about time program and synchronized clock.  RTS at 1428 GMT.

7/8/05

PAD

Computer clock is 7 seconds slow.  Started about time program and synchronized clock.  RTS at 1556 GMT.

7/5/05

PAD

Computer clock is 41 seconds slow.  Started about time program at 1459 GMT.  Clock synchronized by 41477ms.  RTS at 1514 GMT.

6/28/05

CAM

Albert Mendoza notified Site Ops that the MPLHR was experiencing a laser double pulse problem, beginning around May 13.   The laser energy has also dropped from 8W to 6.5W in the last couple of days.  Albert requested Site Ops assistance in making adjustments while he monitored the status via Radmin.  Placed MPL program in ‘Sample Without Saving’ mode at 1944 GMT and changed sample rate to 3 seconds.  Made several incremental adjustments to the Coherent power supply temperature and current settings while observing the effects on the double pulse and laser energy.  Ultimately changed the temperature from 35C to 28.4C and the current from 1200mA to 1330mA.  Returned the sample rate to 30 seconds and turned the save function back on and RTS at 2125 GMT.  Albert will continue to monitor status and advise if further changes are needed.

6/22/05

PAD

Computer clock was 7 seconds slow.  Started about time program and clock adjusted 7.1 seconds.  RTS at 1456 GMT.

6/21/05

TMG

Stopped MPL program @ 1645 GMT and installed RAdmin. Rebooted computer @ 1703 GMT and RTS @ 1714 GMT.

6/21/05

PAD

SDS is unable to collect data from computer.  Rebooted computer at 1423 GMT.  Communication is now restored with SDS.  RTS at 1438 GMT.

6/15/05

PAD

Computer clock is 8 seconds slow.  Started about time program.  Clock adjusted by 7.8 seconds.  RTS at 1440 GMT.

6/14/05

MR

The SDS is not able to collect from the computer.  Rebooted the computer at 1552.  Communications was restored.  RTS at 1607.

6/13/05

PAD

On 6/10/05 at 1400 GMT, observed condensation inside corrector plate.  Removed plate to let instrument dry out over weekend.  At 1837 GMT, Mike Rainwater and I removed bottom plate from instrument and checked optical lens for water spots as directed by mentor.  No water spots were found on lens.  Assembled instrument and restarted.  Notified instrument mentor.  RTS at 1857 GMT.

6/10/05

PAD

On 6/9/05, rainstorm was in area.  MPL enclosure window had leaked water on corrector plate.  Corrector plate was removed and cleaned and no water was observed inside instrument.  On 6/10/05 at 1400 GMT, condensation was observed inside corrector plate.  Stopped program and shut down laser.  Notified instrument mentor.  At 2055 GMT, Mike Rainwater and I removed side panel from laser base and checked for water.  No water was found.  Removed corrector plate and tipped laser.  About a teaspoon of water ran out from behind bottom mirror.  Contacted instrument mentor.  Left telescope open over weekend to dry out.  RTS at 2115 GMT.

6/10/05

PAD

MPL program had locked at 1040 GMT.  Computer is not responsive.  Rebooted computer and MPL program started correctly.  RTS at 1415 GMT.

6/9/05

PAD

Rainstorms at CF with strong winds had forced water inside instrument enclosure.  At 1245 GMT, visiting scientist notified water was leaking onto instrument.  John Schatz placed lens cover on instrument at 1245 GMT.  At 1406 GMT, I shutdown laser and removed corrector plate to clean dust and water spots.  No water observed inside laser housing.  Facilities department removed instrument shelter window and installed window with new caulk.  Replaced corrector plate and restarted laser and computer program at 1606 GMT.  RTS at 1621 GMT.

4/12/05

PAD

MPL program had locked at 0640 GMT.  Computer clock was correct but computer was not responsive.  Laser power supplies are on and laser is working. Rebooted computer at 1427 GMT and started about time program.  Made sure MPL program restarted correctly.  RTS at 1442 GMT.

3/2/05

PAD

Error displayed:  Controller not responding or laser is off.   Coherent power supply shows 0 mA.  Restarted laser at 1450 GMT.  RTS at 1505 GMT.  Assisted facilities department in replacing batteries in UPS.  Instrument operation was not affected during this time.

2/25/05

CAM

SDS collections failed again at 2120 GMT after the system was previously restarted at 2019 GMT.  Check of instrument found everything appeared to be operating normally and plotting data.  Stopped the MPL program and rebooted the laptop at 2132 GMT.  Restarted the laser and RTS at 2137 GMT.

2/25/05

CAM

SDS status page indicated no files had been received from the MPL since 1720 GMT.  Check of instrument found error displayed on computer: Controller Not Responding or Laser is Off.  Computer shows the last average started was at 1601 GMT.  Coherent power supply current reads 0000 and laser is off.  Shut the MPL program down and restarted the laser.  Restarted MPL program and RTS status O.K. at 2019 GMT.

2/25/05

PAD

MPL computer clock is 12 seconds slow.  Started about time program and synchronized clock.  RTS at 1509 GMT.

2/23/05

CAM

Albert Mendoza advised that the energy monitor readings on the MPL were pegged at maximum (7uJ) and requested Site Ops assistance troubleshooting.  Stopped the MPL program at 1940 GMT and placed in test mode with 2 second sample interval.  Reduced the current on the Coherent power supply from its set point of 1200mA and observed that the energy reading eventually decreased as expected with less current.   Slowly increased the current and found that the energy reached 7uJ at around 1128mA.  The laser puts out slightly higher energy with the Coherent power supply than it did with the old Spectra Physics power supply.  The energy monitor had not previously been calibrated to read higher than 7uJ.  Modified the energy monitor calibration file (energy.map) so that it can measure and report higher energy levels.  Returned the Coherent current reading to 1200 and restarted the system in normal sampling mode at 2220 GMT.  RTS status O.K. with energy monitor reading between 7 and 8.

2/23/05

PAD

Error displayed:  Controller not responding or laser is off.  Coherent power supply has power but shows 0mA.  Last successful collection at 1020 GMT. Restarted laser.  RTS at 1539 GMT.

2/22/05

PAD

MPL has error message:  Controller not responding or laser is off.  Computer clock is correct and display shows last average started at 0422:17 GMT.  Coherent power supply has power and shows 0 mA.  Physics keypad is in status mode and shows .04W.  Emission and pulse lights are on.  Spectra physics power supply is on.  Stopped MPL program.  Restarted coherent power supply.  Started MPL program.   At 1557 GMT, display shows cloud height of 1.274 Km. RTS at 1608 GMT.

2/18/05

PAD

Computer clock is 8 seconds slow.  Started about time program.  Program synchronized clock at 1524 GMT.  RTS at 1539 GMT.

1/27/05

PAD

Computer program locked at 1648:36 GMT.  Computer is not responsive.  Power cycled computer and made sure program restarted correctly.  RTS at 2154 GMT.

1/24/05

PAD

Last successful collection was on 1/23/05 at 1120 GMT.  Program is running and seems normal.  Called Ron Reed who forced a collection.  Collection report says no files to get.  Rebooted computer and made sure program restarted correctly.  RTS at 1501 GMT.

1/18/05

PAD

MPLHR program had stopped on 1/18/05 at 0330 GMT. Computer is not responsive.  Power cycled computer at 1449 GMT.  Program restarted correctly.  RTS at 1504 GMT.

1/11/05

PAD

Last successful collection was at 0220 GMT.  Everything on computer display looks normal.  Ron Reed forced collection.  Collection log reports no files kept. Rebooted computer at 1441 GMT.  Checked that SDS had collected data.  RTS at 1456 GMT.

1/3/05

PAD

Program had locked on 12/31/04 at 1648:56 GMT.  Power supplies and laser were operating.  Display shows cloud height of 9.309 Km. Computer is not responsive.    Power cycled computer.  Restarted MPL program.  Display shows cloud height of .225 Km at 1451 GMT.  RTS at 1601 GMT.

12/21/04

PAD

Correction plate window on MPLHR needs cleaned.  Particles cannot be removed with air.  Mike Rainwater and I shut down MPLHR at 2028 GMT.  Removed correction plate window and cleaned with distilled water and isopropyl alcohol.  Installed correction plate and restarted laser.  RTS at 2058 GMT.

12/16/04

PAD

MPL power supply and laser have been frequently shutting down.  Instrument mentor requested UPS checked.  Unplugged UPS at 1957 GMT.  Spectra Physics power supply, Coherent power supply and controller all shut down.  Plugged power supplies, computer and controller into another UPS.  Restarted laser and program.  RTS at 2017 GMT.

12/16/04

PAD

Program and laser was restarted at 1519 GMT.  SDS has not collected since 0820 GMT.  Rebooted computer and restarted MPL program at 1831 GMT.  SDS collected data.  RTS at 1846 GMT.

12/16/04

PAD

Program had stopped at 0644 GMT.  Display shows error:  Controller not responding or laser is off.  Physics keypad shows 0 watts. emission light is on, pulse light is off.  Coherent power supply shows 0 mA.   Power cycled Spectra Physics power supply.  Restarted laser and MPL program at 1519 GMT.   Tried to contact instrument mentor, long distance lines were not working at CF.  E-mailed mentor.  RTS at 1534 GMT.

12/14/04

PAD

At 1517 GMT, display shows controller not responding or laser is off.  Program stopped at 1321 GMT.  Emission light is lit on physics keypad.  Coherent power supply shows 0 mA.  Program had stopped yesterday at 1310 GMT.  Same error message was displayed.  Restarted laser and program at 1517 GMT.  Made sure laser and program restarted correctly.  RTS at 1532 GMT.

12/13/04

PAD

MPL program had stopped at 1310 GMT.  Display shows controller not responding or laser is off.  Keypad shows emission and Interlock.  Coherent power supply shows 0 mA.  Power cycled Spectra physics power supply.  Restarted laser and MPL program.  Made sure program and laser restarted correctly.  RTS at 1525 GMT.

12/10/04

PAD

At 1503 GMT, MPL program was observed running.  Emission light was off on Specta Physics power supply.  Keypad shows TEST.  Coherent power supply shows 0 mA.  Computer clock is 8 sec. slow.  Started about time program and synchronized clock.  Power cycled power supply.  Restarted Laser and program.  Program restarted correctly.  RTS at 1518 GMT.

12/9/04

MR

The program has been shut down due to no response from the laser.  Rebooted the system at 1602 GMT.  System came back up and started taking data at 1615 GMT.  RTS at 1615 GMT.

12/1/04

PAD

Computer had locked at 0047 GMT.  Laser was on but computer was not responsive.  Power cycled computer and made sure program restarted correctly.  RTS at 1517 GMT.

11/22/04

PAD

Last successful collection was on 11/20/04 at 0120 GMT.  Display shows Controller not responding or laser is off.  Coherent power supply shows 0 mA.   Restarted Laser and MPL program at 1450 GMT.  Program restarted correctly.  At 1455 GMT, display shows cloud height of 1.154 Km, which is correct for conditions.  I was notified at 1700 GMT that SDS system had not collected data from instrument.  Stopped program and rebooted computer at 1715 GMT.  Made sure program restarted correctly and data was successfully collected at 1820 GMT.  RTS at 1730 GMT.

11/8/04

PAD

At 1525 GMT, observed computer program had stopped at 1504:36 GMT.  Computer was not responsive.  Laser was on.  Power cycled computer at 1529 GMT.  Started about time program and made sure program restarted correctly.  At 1533 GMT, display shows cloud height of 6.5 Km. RTS at 1540 GMT.

11/1/04

PAD

MPL computer clock is slow.  Synchronized clock at 1450 GMT.  About time shows corrected error of 5650 ms.  RTS at 1505 GMT.

11/1/04

PAD

MPL last successful collection was on 10/29/04 at 2320 GMT.  Error message shows controller not responding or laser is off.  Physics keypad status shows 6 on display, emission light and pulse light is on.  Keypad is not responsive.  Power cycled Spectra Physics laser power supply at 1447 GMT.  Restarted laser and MPL program.  Made sure program restarted correctly.  RTS at 1502 GMT.

10/28/04

PAD

MPL last successful collection was at 0120 GMT.  Storms moved through area at that time.  Display shows Controller not responding or laser is off.  Coherent power supply had power but laser was off.  Restarted laser and MPL program at 1343 GMT.  Made sure instrument restarted correctly.  RTS at 1358 GMT.

10/26/04

PAD

At 1542 GMT, observed MPL program had stopped, laser was off and keypad did not show any readings and was unresponsive.  Site is experiencing power bumps at this time.  Power cycled Spectra Physics laser power supply and restarted laser and MPL program.  At 1548 GMT, program had restarted correctly and shows a cloud height of 4.542 Km. RTS at 1557 GMT.

10/21/04

PAD

Computer clock is consistently 2 sec. slow.  SDS computer is sending hourly notices to Ron Reed that clock needs adjusting.  Computer is running About time 4.8 software.  Adjusted time at 1511 GMT.  Clock adjusted .540ms.  Changed settings from updating time every 15 min. to every 5 min. as per Ron Reed's request.  RTS at 1526 GMT.

10/20/04

PAD

At 1347 GMT on 10/19/04, MPL program had stopped.  Coherent laser diode had power but laser was not on.    Restarted laser and MPL program at 1353 GMT.  Made sure program restarted correctly.  RTS at 1408 GMT.

10/14/04

PAD

MPL last successful collection was on 10/13/04 at 1820 GMT.  Computer was running, time was correct and program seems to be running normal.  Power cycled computer at 1418 GMT to reestablish communications.  RTS at 1433 GMT.

9/10/04

MR

The instrument mentor called and wants some system settings changed.  Called the instrument mentor back.  He walked me through the procedure to check and change the output power settings for the MPLHR.  RTS at 2230 GMT.

8/18/04

PAD

While performing PM rounds on MPL instrument, I accidentally stopped emission at 1404 GMT.   Restarted emission and program at 1404 GMT.  RTS at 1419 GMT.

8/13/04

PAD

Computer clock is 8 sec. slow.  About time icon is not displayed on taskbar.  At 1456 GMT, started about time program and synchronized clock.  RTS at 1511 GMT.

8/12/04

MR

The instrument mentor has sent a replacement keypad and laser power supply to replace the one currently being used in the system.  R/R the laser power supply and keypad.  Called the mentor and he had me double check everything and then powered the unit up.  Everything checked ok.  RTS at 1630 GMT.

8/10/04

PAD

Computer is 10 sec. slow.  About time Icon is not on taskbar.  Started about time program at 1835 GMT.  RTS at 1850 GMT.

8/10/04

PAD

MPL program stopped at 1645 GMT on 8/9/04.  Last successful collection was at 1820 GMT.  Computer displayed error:  Controller not responding, or Laser is off.  Sampling process has stopped.  Emission and pulse lights on physics keypad were off.  Restarted emission and pulse on physics keypad and MPL program.  Made sure instrument restarted correctly.  RTS at 1401 GMT.

8/9/04

PAD

Last successful collection was on 8/7/04 at 1520 GMT.  Error displayed:  Controller not responding, or Laser is off.  Sampling process has stopped.  Program had stopped at 1325 GMT.  Checked all connections between Laser diode, physics keypad, computer, and controller/ scaler.  Power cycled Laser diode module and rebooted computer.  Restarted MPL program.  RTS at 1419 GMT.

8/5/04

PAD

MPLHR program stopped at 0123 GMT.  Emission light was off on physics keypad.  Started emission on physics keypad and restarted MPL program at 1449 GMT.  RTS at 1504 GMT. 

8/4/04

PAD

Last successful collection was at 0220 GMT.  Error displayed:  Controller not responding or laser is off.  Checked connections at physics keypad, laser diode module, and computer.  Started pulse and emission at physics keypad at 1352 GMT.  Made sure program restarted correctly.  RTS at 1405 GMT.

8/2/04

PAD

Last successful collection was on 7/30/04 at 2320 GMT.  At 1412 GMT, computer displayed error: Controller not responding or Laser is off.  Emission and pulse light was off.  Power cycled Spectra Physics laser diode module at 1415 GMT.  Restarted MPL program.  RTS at 1430 GMT.

7/28/04

MR

The site power will be going down and this instrument needs to be shut down.  Shut the instrument down between 1439 and 1611 GMT.  RTS at 1612 GMT.

7/27/04

TMG

IAW BCR 796: Uninstalled D4 software @ 1525 GMT and installed AboutTime @ 1540 GMT.  Performed a set time and the clock was adjusted +7178 ms. RTS 1541 GMT.  Note: AboutTime options were set to check time every 15 minutes with 198.124.99.74 timeserver and to start program at startup.

7/26/04

MR

The instrument has missed the last couple of collections.  Rebooted the computer at 1548 GMT to reestablish the communications with the SDS.  RTS at 1603 GMT.

7/26/04

PAD

Computer clock is 23 sec. slow.  Synchronized computer clock at 1428 GMT.  RTS at 1443 GMT.          

7/26/04

PAD

Error message displayed:  Controller not responding or Laser is off.  Emission light on physics keypad was off.

7/23/04

PAD

Coherent laser diode supply was installed previously today.  Instrument mentor requested files moved and instrument restarted.  At 1754 GMT, instrument mentor called.  Copied Coherent. trm file to C/Current/Run Me/ Run_Last folder as per mentor's request.  Powered up MPL scaler and control.  Made sure program restarted correctly.  RTS at 1824 GMT.

7/23/04

MR

The new MPLHR power supply has come in and needs to be installed.  Installed the MPLHR power supply as per Conner Flynn’s written instructions.  Restarted the MPL program and had a "HW controller not responding error"  Notified Conner Flynn by voice mail and will await his instructions.  RTS at 1600 GMT.

7/6/04

PAD

Computer clock is 41.78 sec. slow.  Synchronized clock at 1534 GMT.  RTS at 1549 GMT.

6/30/04

MR

The power to the site was lost at 1330 GMT.  While power was out, the system was physically turned off so a controlled power up could be performed.  Power was restored to the site at 1455.  Powered up the system at 1553 GMT.  System came up and rebooted ok.  RTS at 1608 GMT.

5/21/04

PAD

Computer clock time is 33 sec. slow.  Synchronized clock at 1726 GMT.  RTS at 1741 GMT.

4/12/04

PAD

After cleaning window, computer displayed an error box:  A required resource was OK.  Connection between laptop and instrument is extremely sensitive to movement.  Connection was disrupted.  Restarted MPL program at 1548 GMT.  Made sure program started correctly.  RTS at 1603 GMT.

2/5/04

PAD

Computer program had stopped at 1547 GMT on 2/4/04.  Restarted computer program at 1622 GMT.  Made sure program started correctly.  RTS at 1629 GMT.

12/15/03

PAD

During PM rounds, after checking SD and LO values, computer showed following message:  A required resource was.  MPL program stopped.    I clicked OK on the message, MPL program did not restart.  Closed and restarted MPL program at 1545 GMT.  Made sure MPL program restarted correctly.  RTS at 1600 GMT.

10/9/03

PAD

MPLHR last successful collection of data was on 10/9/03 at 1720 GMT.  At 2107 GMT, the program seemed to be running normally except for the 30 second countdown clock.  Stopped MPL program at 2107 GMT and restarted the program at 2108.  MPL program seemed to operate normally.  RTS at 2122 GMT.

9/11/03

KC

The program and computer clock are running 45 seconds behind the GMT clock. Reset the program and computer at 1508 GMT. RTS at 1515 GMT.

9/9/03

PAD

Computer clock was 2 min 50 sec. slow.  Adjusted computer clock at 1504 GMT.  RTS at 1514 GMT.

8/22/03

PAD

Computer locked up while checking SD 255 values.  Rebooted computer at 1452 GMT. Checked display for accurate data display and graphing.  RTS at 1600 GMT.

5/17/03

CAM

The DS View instrument status page indicated no files had been received from the MPL since 0300 GMT.  Check of instrument found that it was running, but just not sending files.  Rebooted the MPL laptop at 1815 GMT.  RTS at 1817 GMT.  Note the optical trailer UPS experienced a power outage from 0345 GMT to 1746 GMT, which caused the network hub to go down.  All data files were recovered when the UPS power, and thus network connection, was restored.

5/15/03

JEM

A fatal exception error has occurred.  The error is: "An exception OE has occurred at 0028:C001545A in VxD - - -.  Pressed the spacebar to clear the error message.  The MPL program was still running in the background and was saving data files.  A current cloud height of 5.161 was observed which appears to agree with current conditions.  RTS @1842 GMT.

5/7/03

JEM

The MPL laptop was observed locked up displaying a time and date of 5/7/03 at 0713 GMT.  The MPL laptop was unresponsive to keyboard and mouse commands.  Powered off the MPL computer using the power button @1402 GMT.  Powered the laptop back on @1403 GMT.  The MPL was back online and operating normally @1407 GMT.  No cloud height was displayed and this agreed with the current conditions.  RTS @1416 GMT.

5/5/03

JEM

The MPL has not collected since the 1720 GMT collection on 5/2/03. The MPL program was observed running as expected, but no data files were being created.  Power cycled the MPL laptop @1509 GMT.    At 1515 GMT the MPL was observed operating normally and saving data files.  RTS @1520 GMT.

5/5/03

JEM

The MPL laptop time is off and needs to be set.  Set the laptop time @1507 GMT using the D4 program.  The laptop time was slow 18min 22sec.  RTS @1520 GMT.

5/2/03

MR

The system clock has drifted outside acceptable limits.  Reset the system clock at 1625 GMT.  RTS at 1640 GMT.

4/25/03

JEM

The SDS has not been collecting files from the MPL.  The MPL appeared to be operating normally and was generating new files.  Power cycled the MPL @1410 GMT.  At 1415 the MPL was back online and operating normally.  RTS @1422 GMT.

4/18/03

JEM

The MPL time is off by 00:03:12.  Adjusted the MPL clock @1642 GMT.  RTS @1645 GMT.

4/16/03

JEM

No beam output was observed and the following error was observed.  "Controller not responding or laser is off.  Sampling process stopped".  It appears that during the power outage the laser lost is setting of SD = 15 (on) and is now reading SD = 0 (off).  Using hyper terminal, typed in SD=15 @1416 GMT to turn the laser back on.  The laser is now on and a beam output is observed.  Restarted the program @1418 GMT.  At 1421 GMT the MPL was operating normally displaying a cloud height of 1.634km.  RTS @1430 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage.  Power  was restored to the Central Facility at approximately 1945 GMT.  The AOS was brought back online @2010 GMT.  RTS @2015 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage @1845 GMT.  Power was restored to the Central Facility at approximately 1945 GMT.  The MPL was brought back online 1950 GMT.  Set the Diode Current value to 15 and set the laser on/off value to 3 @1955 GMT.  RTS @2000 GMT.

1/30/03

JEM

Albert Mendoza is on site today making necessary repairs to the MPLHR.

1/15/03

JEM

The MPL graph is blank, and the background is reading 0.0 e+0.  Notified Connor Flynn.  He advised that the detector is bad and that he and Albert Mendoza will have to come out and replace the MPL.  Shut the MPL down  @1855 GMT.  The MPL will remain shutdown until it is replaced.  Did Not RTS.

1/8/03

MR

The instrument mentor has requested that the laser power supply be changed out.  Changed out the laser *power supply.  Called the mentor and he had me perform some performance tests and concluded that the laser may need an alignment to completely cure the low power problems.  He will issue a work order when he is ready for the alignment.  RTS at 2000.   *Old – 3689   *New - U3680 

1/7/03

JEM

The following error was observed on the laptop.  “An exception OE has occurred at 0028:C001545A in VxD ---.”  Pressed enter to clear the error.  Shutdown the computer @1640 GMT.  The computer became hung up for several minutes while shutting down.  Pressed and held the power button until the laptop shut off @1644 GMT.  Powered the laptop back on @1645 GMT.  At 1650 GMT the MPL was observed operating normally.  RTS @1652 GMT.

12/31/02

JEM

The MPL computer locked up this morning @0844 GMT.  The laptop was unresponsive to any commands.  Held the power button down for about 10 seconds to turn the laptop off @1531 GMT.  Powered the laptop back on @1532 GMT.  After the computer booted and the program started, the MPL was observed operating normally @1540 GMT.  RTS @1545 GMT.

12/16/02

JEM

The time is wrong on the MPL computer.  Synchronized the time using the D4 program @1556 GMT.  The time was off by -00:10:05.  RTS @1611 GMT.

12/5/02

MR

The SDS was not able to communicate with the laptop.  Rebooted the laptop computer.  RTS at 1557 GMT.

11/19/02

MR

The computer had a fatal exception error displayed and was locked up.  Rebooted the computer at 1944 GMT.  System came back up and started working at 1949.  RTS at 2000 GMT.

10/22/02

MR

The program seemed to be locked up.  The last know data point on the graph was at 2023 GMT on 10/21/02.   Rebooted the computer at 1629 GMT.  The system came back up and started collecting data at 1633 GMT.  RTS at 1645 GMT.

10/17/02 – 10/18/02

C. Flynn

The MPL was taken off-line at 1428 GMT on 10/17 for the following maintenance:

The MPL optics were cleaned and realigned.  Substantial improvements in

signal to noise were yielded (about a factor of six increase).  Conducted overlap measurements, adjusted the optics collimation and provided training for such.  Collected the following calibrations measurements:

1. Darkcounts measurement 02101817.11W

2. Afterpulse measurement 02101817.36W

3. Overlap determination 02101820.37W

4. Energy monitor calibration energy.20021018.map

These files have been provided to James Campbell of NASA GSFC for generation of correction files for MPL_Nor.  The MPL was returned to operation on Friday Oct. 18, 2200 GMT and appears to be in good working order except that currently the files are not being collected properly due to an unrelated issue with the data system.

9/26/02

JEM

The MPL has been shut down since 9/24/02 so it could be dried out.  The MPL is dried out and needs to be brought back online.  Powered on the scalar controller @1515, the laser @1516 and the laptop @1517 GMT.  After the laptop completed booting up everything appeared normal.  No clouds were being indicated which agreed with the current conditions.  Status OK.  RTS @1525 GMT.

9/24/02

JEM

The MPL had condensation on the corrector plate again.  Powered the MPL down @1610 GMT.  Removed and dried the corrector plate.  Turned the heater in the room to 80 degrees.  Will leave the system open with the heater up until 9/26/02 to completely dry any moisture from the inside of the MPL.  RTS @1625 GMT.

9/23/02

JEM

Condensation has been observed on the bottom of the corrector plate since the window was blown off last weekend.  The mentor has several things that need to be checked.  Powered off the MPL @1445 GMT.  Took the telescope off of the box and dried all traces of water from within it.  Removed 3 sides of the MPL base.  Checked inside the base for water, or traces of water on the optics.  No evidence of water was found anywhere inside the base of the MPL.  Used a heat gun to ensure that the telescope was totally dry.  Used methanol to clean the 2 mirrors and the corrector plate.  Took several pictures of the MPL during this process to send to the mentor.  Re-assembled the MPL and powered up the system @1610 GMT.  Status OK  RTS @1615 GMT.

9/20/02

JEM

Condensation was observed inside the corrector plate.  Removed the corrector plate @1415 GMT and cleaned the condensation from it.  Lined up the alignment marks on the correct plate and reinstalled it @1425 GMT.  RTS @1426 GMT.

9/17/02

JEM

The MPL has not collected or ingested since 1702 GMT on 9/16/02.  Was informed by Tim Grove that the MPL Ingest program no longer needs to run because it FTP's files to the SDS.  Closed MPL Ingest and removed it from the startup group @1340 GMT.  Status OK.  RTS @1355 GMT.

9/16/02

JEM

The window has been replaced and the system needs to be powered on.  Powered the system on @1743 GMT.  After the system had powered up, set the output power to 1.00 watts and verified that the pulse and emissions lights were on.  The computer displays an error message stating "HW: Laser is not running".  Checked all connections and settings.  The cable running from the laptop to the laser was slightly unplugged.  Plugged the cable in and restarted the MPL software.  Status OK.  RTS @1758 GMT.

9/16/02

JEM

The port window has been blown off the building.  The MPL got significant amounts of water inside due to the window being blown off.  Totally powered off the system and removed the water from the telescope. Will wait for the window to be replaced by facilities before powering up the system.  RTS @1425 GMT.

7/19/02

KT

Instrument mentor requested the change out of the *Laser Supply fan.  At 1439 GMT stopped the program and powered down the system.  At 1449 GMT removed the Laser Supply unit.  At 1523 GMT removed and replaced the fan.  At 1540 GMT put the Laser Supply unit back together and installed into instrument rack.  At 1600 GMT after connecting the cables to the unit, power was turned on.  At 1609 GMT the system came on line but the beam was not visible.  After checking cable connections and power out put to the laser, the system was rebooted.  At 1620 GMT when system came on line the beam was still not visible.  At this time the system was powered off again and Albert Mendoza was notified.  After talking with Albert, I went back through the connections and found that the fiber cable was not seated all the way possibly causing the beam not to transmit at full strength.  At 1700 the system was powered on.  After the MPLHR came on line, the beam was transmitting at max and the system was operating within range.  RTS at 1705 GMT.  Status OK.  *Old - 6213E  *New - 9350

7/9/02

JS

The PC ingest was turned off at SGP on 7/9/02. The SGP is processing and shipping MPL data.

7/9/02

CT

The Ingest program appeared to have terminated. Message displayed was Path/File error. The computer was rebooted at 1437 GMT. The mentor has been notified of this reoccurring problem. RTS at 1452 GMT.

7/8/02

CT

A Path/File Error was displayed. The Ingest appeared to have terminated. The computer was rebooted at 1452 GMT. RTS at 1507 GMT.

7/5/02

CT

Message was displayed that a "Path/ File access error" had occurred. The Hypersnap program was terminated. The computer was rebooted at 1450 GMT. RTS at 1505 GMT.

7/3/02

CT

Message displayed " Path/File error" This message did not indicate the Ingest was affected. But the Ingest appeared to be hung up. The computer was rebooted at 1441 GMT. Status OK. RTS at 1456 GMT.

7/2/02

CT

An error message was present on the computer display. The message was "Path/File access error"The computer was rebooted at 1444 GMT. Status OK. RTS at 1459 GMT.

7/2/02

CT

The computer clock was approximately 3 minutes slow. The computer clock was adjusted at 1446 GMT. RTS at 1501 GMT.

7/1/02

KT

There was a Run Time error 75 message that appeared.  The ingest needed to be rebooted.  At 1512 GMT closed the ingest window.  At 1514 GMT rebooted the ingest.  Status ok.  RTS at 1525 GMT. 

6/28/02

CT

The Ingest program was terminated. The Ingest program was started at 1525 GMT. RTS at 1540 GMT.

6/27/02

MR

The ingest was observed to have stopped with an runtime error 75.  Restarted the ingest.   RTS at 1456 GMT.

6/24/02

KT

The ingest had quit running.    At 1400 GMT powered off system completely.  At 1403 GMT powered system on.  At 1405 GMT system came on line and ingest was running normally.  RTS at 1413 GMT. 

6/21/02

CT

The Ingest program was terminated. The Ingest was started at 1426 GMT. Status OK. RTS at 1441 GMT.

6/20/02

KT

There were no files ingested.  At 1853 GMT found that the ingest had stopped.  Rebooted the ingest.  RTS at 1908 GMT.

6/19/02

MR

The ingest has stopped.  Restarted the ingest at 1445 GMT.  RTS at 1500 GMT.

6/17/02

CT

The Ingest program was terminated. Started the Ingest program at 1459 GMT. RTS at 1514 GMT.

6/14/02

MR

The system lost power on 6/13/02 due to power outage to the site.  The instrument power switch was turned off until power was restored to the site.  Power has been restored to the site.  Powered on the instrument at 1440 GMT.  System came back up and started taking data at 1443 GMT.  RTS at 1455 GMT.

6/13/02

MR

The site has lost power due to storms in the area at 1415 GMT.  The instrument is shut down at this time.

6/10/02

CT

The Ingest program was terminated. "Run-Time" error and "Path/File Access" error was displayed. The Ingest program was restarted at 1453 GMT. RTS at 1508 GMT.

6/7/02

CT

The Ingest was terminated. "Run-Time Error '75' " and "Path/File Access Error" was displayed. The Ingest was started at 1415 GMT. Status OK. RTS at 1430 GMT.

6/3/02

TMG

The MWR-HR has not collected any data since 5/31/2002 at 20:03 GMT. T/S indicates that the MPL menu states that the MPL is in an idle state. Rebooted computer at 1620 GMT. MPL menu displayed sampling 02060316.25W. System RTS @1635 GMT.

5/31/02

KT

The corrector plate had two dried water spots and needed to be cleaned.  At 2022 GMT powered down the laser by using the controller.  Cleaned the dried water spots off the corrector plate.  At 2026 GMT powered the laser up.  RTS at 2035 GMT. 

5/31/02

CT

The corrector plate appeared to have two dried water spots on it. At 1445 GMT a message was left on the mentor's voice mail concerning the situation. Waiting on the mentor to advise on action to be taken. RTS at 1500 GMT.

5/23/02

TMG

2030 GMT to 2035 GMT.  D4 program was set to adjust the computer clock by the old ntphost which is no longer available. Set D4 program to adjust the computer clock to the new ntphost 99.74 and the program adjusted the time +10.29 sec. Note: The D4 program will set the computer clock once every 15 minutes. RTS. 

5/23/02

CT

The computer clock was 5 minutes off. The computer was rebooted at 1359 GMT. The computer clock returned to the correct time. Status OK. RTS at 1414 GMT.

5/22/02

CT

The Ingest program was terminated. The message displayed was "Run time Error 75" and Path/File Access Error" The Ingest program was restarted at 1510 GMT. RTS at 1525 GMT.

5/21/02

CT

The Ingest program had terminated. The message displayed was "Run Time Error 75" And  "Path/File Access Error".  The Ingest program was restarted at 1440 GMT. RTS at 1455 GMT.

5/20/02

CF

The Ingest program was terminated. Message stated "Run time error 75". Rebooted the program at 1448 GMT. RTS at 1503 GMT.

5/9/02

KT

After bringing the instrument online, the data graphs were scrolling but no data was being displayed.  At 2103 GMT called Albert Mendoza and informed him of the situation.  After trouble shooting the instrument, the detector was found to be the suspected root cause of the problem.  At 2109 GMT powered the instrument down until further notice.  RTS at 2115 GMT.

5/9/02

CT

During inspection there was no laser emission output.  At 1533 GMT rebooted the laptop.  After rebooting, there was still no emission output.  At 1540 GMT observed that the emission light was not lit on the hand controller.  At 1542 GMT pressed the emission button on the controller and the beam was visible and operating.  RTS at 1546 GMT.

4/15/02

CT

The computer had experienced an Exception 0E Error. Power cycled the computer at 1508 GMT. This did not bring the computer up. The battery pack was removed then replaced. Status OK. RTS at 1533 GMT.

4/10/02

SS

The instrument missed several collections.  The instrument computer was checked and indicated an error 75 on the data ingest program.  The ingest program was restarted at 1430 GMT.  When the restart was complete, the instrument collections returned.  RTS 1505 GMT.  Status was OK at this time.

4/9/02

SS

The MPL was not collecting data.  The MPL data ingest program was showing an error.  The ingest program was restarted at 1640 GMT.  When the restart was complete, the data collection resumed as normal.  RTS 1655 GMT.  Status was OK at this time.

4/8/02

MR

The computer locked up while POP4 was rebooting.    Rebooted the computer at 1503 GMT.  RTS at 1518.

4/1/02

CT

There was a fatal exception error stating that 0E has occurred at 0028: C14BC80E in VXD VIDI (01) +0000024E. The current application will be terminated. Rebooted the computer at 1607 GMT. Status OK. RTS at 1622 GMT.

3/21/02

MR

The communications with the instrument have failed.  The computer had a fatal exception error displayed on the screen.  I clicked ok and the error closed out and the program restarted.  I rebooted the computer at 1529 so that the ingest would restart and a new connection with the SDS would be established.  The instrument restarted at 1533 GMT.  RTS at 1543 GMT.

3/11/02

MR

The power to the site was lost over the weekend.  The instrument was found to be off.  Powered the system on at 1629 GMT.  System came up and started taking data at 1632 GMT.  RTS at 1644 GMT.

3/8/02

CT

The Ingest program was terminated. "Run time error 75" and "Path/file access error" was displayed. Restarted the Ingest program at 1449 GMT. RTS at 1504 GMT.

3/7/02

KT

The Ingest program had terminated. The message displayed was "Ingest run time error 75"Reset the Ingest at 1533 GMT. RTS at 1548 GMT.

3/5/02

CT

The computer clock was not correct. Reset the computer clock at 1550 GMT. RTS 1605 GMT.

3/5/02

CT

Two Ingest programs appeared to be running at the same time. Shut down the Ingest programs and restarted the Ingest program at 1550 GMT. RTS at 1605 GMT.

3/4/02

CT

The Emission light was not on. Turned the Emission light on at 1608 GMT. RTS at 1623 GMT.

3/4/02

CT

There was a error message displayed "controller not responding".  Power cycled the computer at 1607 GMT. RTS at 1622 GMT.

2/25/02

CT

The Ingest program had terminated. Restarted the Ingest program at 1514 GMT. RTS at 1529 GMT.

2/22/02

CT

The clock on computer display was not correct. Reset the computer time display. RTS at 1705 GMT.

2/20/02

MR

The instrument mentor wants the instrument checked for some possible alignment and power problems.  Called Conner Flynn and he remotely logged on and checked a few things.  He had me check for output and the output looked fine.  He had me check the emission value and it was set at 1.01 as it should be.  He concluded that the detector may have been knocked out of alignment and the instrument will need to have an alignment performed on it.  He said he would arrange this and would let us know what and when it will happen.  RTS at 2045 GMT.

2/19/02

CT

The Ingest program had terminated. Restarted the Ingest program at 1608 GMT. RTS at 1623 GMT.

2/15/02

CT

The Ingest program had terminated. Restarted the Ingest program at 1642 GMT. RTS at 1657 GMT.

2/11/02

MR

The AC power to the site was lost on 1/30/02 due to local ice storm.  The instrument laser head had been removed and stored in a safe climate controlled location.    Reinstalled the instrument laser head.  Restarted the computer.  The system came up ok.  Status ok.  RTS at 2215.

12/17/2001

CJ

The Ingest program had terminated. Restarted the Ingest program at 1618 GMT. Status OK. RTS at 1633 GMT.

12/7/01

MR

The instrument mentor has requested that the MPLHR be checked due to he has had no data reported for several days.  T/S found the instrument operating OK but the computer dept. could not ping the instrument.  Rebooted the computer at 2205 GMT.  After the reboot the computer could be pinged remotely.  Called Conner Flynn and he was able to log onto the computer.  RTS at 2215 GMT.

12/03/01

CJ

The Ingest program was terminated. Restarted the Ingest program at 1542 GMT. RTS at 1557 GMT.

11/29/01

CJ

Ingest program had terminated. Restarted Ingest program at 1650 GMT. RTS at 1705 GMT.

11/26/01

CJ

The Ingest program had terminated. Restarted the Ingest program at 1525 GMT. RTS at 1540 GMT.

11/20/01

MR

The power to the site will be shut down for about 1 hour for a site water leak repair.  The instrument needs to be gracefully shut down.  Shut the instrument down at 1444 GMT.  Power to the site was actually down for approximately 1 hour.  Powered up the instrument at 1657.  RTS at 1700.

11/19/01

KT

The ingest program had stopped.  At 1635 GMT closed the ingest window and restarted the program.  Ingest of data files began.  Status ok.  RTS at 1648 GMT. 

11/16/01

KT

The ingest program had stopped and needed to be rebooted.  At 1546 GMT closed the ingest window and restarted the ingest program.  At 1549 GMT the program came on line and began to ingest files.  RTS at 1600 GMT. 

11/13/01

CJ

Ingest program had terminated. Restarted Ingest program at 1537 GMT. RTS at 1552 GMT.

11/09/01

CJ

Error message that read "Illegal operation" at 1527 GMT. Rebooted computer at 1536 GMT. No response. Power cycled computer at 1537 GMT. RTS at 1552 GMT.

11/08/01

CJ

Found the Ingest program had stopped at 1541 GMT. Restarted the Ingest program at 1542 GMT. RTS at 1557.

11/07/01

CJ

Ingest program had stopped at 1518 GMT. Restarted Ingest program at 1518 GMT. RTS at 1533 GMT.

11/05/01

CJ

Ingest program had terminated. Started Ingest program at 1621 GMT. RTS at 1636 GMT.

11/02/01

CJ

Found that the Ingest program had terminated at 1708 GMT. Started the Ingest program at 1710 GMT. RTS at 1725 GMT.

10/30/2001

TMG

The MPL-HR computer has Norton v5.0 antivirus installed and will be needing an up-to-date virus scans. Technician performed a live up to date scan of the system. No viruses were found. Computer RTS at 1802 GMT.

10/30/01

CJ

Found the Ingest program needed to be restarted at 1458 GMT.  Restarted Ingest program at 1459 GMT. RTS at 1514 GMT.

10/29/01

CJ

Found that the ingest needed to be restarted at 1805 GMT. Restarted ingest. RTS at 1820 GMT..

10/26/01

KT

The MPL mentor arrived on site for troubleshooting and repair of the MPL instrument.  Albert Mendoza arrived on site and began troubleshooting of the instrument.  Mr. Mendoza left the CF and was able to bring the instrument back on line.  The *PC was removed and replaced as well as the **Laser Diode Supply.  RTS at 1717 GMT on 10/25/01.  *Old PC: IBM ThinkPad, DOE Property# WD22793  *New PC: IBM ThinkPad A22m, DOE Property# PT05821   **Old Laser Diode Supply: Spectra Physics WD# 27546  **New Laser Diode Supply: Spectra Physics WD# 27542

8/6/01

JEM

At 1430 GMT found the MPL program locked up. Shut down, then restarted at 1431 GMT. Program was operational at 1445 GMT.

8/2/01

KT

Instrument mentor states that the data is corrupted. At 1310 GMT while inspecting the instrument, the plastic cover that is used to cover the beam prior to cleaning was still installed.  At 1313 GMT the cover was removed.  RTS at 1325 GMT. 

8/01/01

JEM

At 1410 GMT while checking data on MPL laptop noticed that ingest program had locked up. The error message; 'run time error 75/ path/file access error. Shut down ingest and restarted at 1410 GMT. Ingest was operational at 1425 GMT.

7/31/01

JEM

At 1430 GMT noticed that the ingested data program had locked up displaying a 'run time error 75; path/access error message.  This occurs almost daily. At 1430 GMT shut down ingest program, then restarted the ingest program.  The instrument was operational at 1445 GMT.

7/26/01

JEM

At 1435 GMT noticed that the ingest program had locked up from early this morning (0001 GMT). Stopped ingest and restarted it.  Returned to service at 1450 GMT.

7/20/01

KT

There was a run time error 75 being displayed. At 1413 GMT closed the ingest window.  At 1415 GMT restarted the MPLHR ingest.  RTS at 1425 GMT. 

7/18/01

JEM

At 1438 GMT noticed that ingest had discontinued, displaying the error message; 'Run time error 75; Path/ file access error. At 1439 GMT shut down data ingest program.  Also, the sampling program was not counting down so a hard reboot (shut system down, then restarted) was performed. System was operational at 1455 GMT.

7/16/01

JEM

At 1440 GMT ingest had discontinued due to a 'Run time error 75'.  This error occurs daily.  There had been no ingested data since 7/15 @ 0027 GMT. At 1440 GMT shut down ingest and restarted .  Ingest was operational at 1455 GMT.

7/12/01

JEM

At 1415 GMT noticed that the data ingest program on the MPLHR laptop computer had locked up, and was giving a ‘Run time error 75’ error message. (This seems to occur daily). I  proceeded to shut down ingest program and restart it.  Ingest data program was operational at 1430 GMT.

7/10/01

JEM

At 1512 GMT while checking data noticed that ingest had shut down giving the error message; 'Run time error 75; Path/access error.  (This is almost a daily occurrence.)  At 1512 GMT shut down MPL data ingest.  Rebooted at 1512 GMT.  Program was operational at 1525 GMT.

7/9/01

JEM

Instrument's ingest display was showing the error message; 'Run time error 75. Path/access error at 1348 GMT.  Shut down ingest at 1348 GMT. Rebooted ingest and program was on line at 1400 GMT.

7/6/01

JEM

At 1440 GMT while checking ingest data there was an error message; 'Run time error 75; Path\access error. (This seems to be daily). At 1441 GMT shut down the ingest program and rebooted it .  Data ingest was operational at 1455 GMT.

7/3/01

JEM

At 1450 GMT noticed a 'run time error 75' on MPLHR laptop. 'path/acess denied' while checking the ingest.  Used shut down procedure to disable ingest. Then performed a soft reboot and ingest was operational by 1505 GMT.

7/02/01

JEM

At 1436 GMT the ingest for the MPL was locked up with an error message that read, ‘run time error 75’ and ‘path/access error’.  After rebooting laptop, the ingest was functioning normally at 1450 GMT.  Seems to be another power supply problem.

6/29/01

MR

The ingest has stopped with a run time error 75. Restarted the ingest from the daesktop.  RTS at 1435.

6/27/01

JEM

The ingest on the mplhr had locked up. The error message; “Run-time error 75” was displayed at 1450 GMT. I tried to reboot ingest program, but another error message; “Path/file access error” was displayed. I turned laptop off, then restarted at 1500 GMT. Both laptop, and ingest program were operational at 1510 GMT.

6/25/01

JEM

The ingest on the mplhr had locked up. The error message; “Run-time error 75” was displayed.  At 1450 GMT the program was shut down; then rebooted. The program is operational at this time.

6/21/01

KT

There was a run time error 75 displayed.  At 1424 GMT closed run time error window.  At 1426 GMT restarted the MPL ingest.  At 1428 GMT ingest began.  RTS at 1437 GMT.

6/14/01

JEM

The MPL laptop ingest program had a ‘run time error 75’ on it at 1345 GMT. Restarted ingest program at 1400 GMT. Program is operational at this time.

6/11/01

JEM

The laptop for the MPLHR instrument had a ‘run time error 75’ displayed while doing PM's. Rebooted laptop and checked ingest. The program recognizes files. Presently operational.

6/04/01

JEM

While checking ingest a "runtime error 75" message appeared at 1359 GMT. Shut down and rebooted laptop, and made sure ingest and other programs were operational at 1415 GMT.

5/30/01

MR

The display on the MPLHR laptop was locked up. Rebooted at 1405 GMT. Ingest and sampling returned. System is now operational.

5/16/01

MR

The ingest has stopped with an "run time error 75". Restarted the ingest at 1505. RTS at 1520.

5/15/01

MR

The ingest has stopped. Rebooted the ingest. RTS at 1532 GMT.

5/8/01

MR

The ingest has stopped with an error 75. Restarted the ingest. RTS at 1445.

5/7/01

MR

The ingest has stopped. Restarted the ingest at 1500 GMT. RTS at 1515 GMT.

5/4/01

MR

The ingest has stopped with an error 75. Restarted the ingest. RTS at 1515 GMT.

5/3/01

CAM

SDS comlog reports indicated that the MPL collection failed at 1904 GMT with the error message No Files To Collect. T/S found that the MPL and ingest programs were both running without errors. The laptop was rebooted at 2005 GMT in an attempt to solve the collection problem. Check of SDS comlog reports later found that the MPL had successfully collected data at 2004 GMT just prior to the reboot, so the reboot was actually not necessary.

5/3/01

CAM

SDS comlog reports indicated that no files had been collected from the MPL since 0004 GMT today with the error message No Files To Collect. T/S found that the ingest program had a runtime 75 error on the laptop. Restarted the ingest program at 1655 GMT. 18 files were found and ingested. The following collection was successful at 1709 GMT.

5/2/01

CAM

SDS comlog reports indicated that no files had been collected from the MPL since 0004 GMT on 4/30. An FTP error was occurring. Attempts to ping the laptop over the network were unsuccessful. Checked the instrument status and found there was no link light on the network adapter. Rebooted the laptop at 1825 GMT and the link light then lit on restart. The laptop could then be pinged over the network. RTS status O.K. at 1830 GMT.

4/27/01

JEM

The MPL has stopped ingesting files again. The MPL had a runtime 75 error again today. Instead of just restarting the ingest again, rebooted the entire MPL computer @1425 GMT. At 1430 the MPL was observed to be functioning properly. RTS @1430 GMT.

4/26/01

JEM

The MPL stopped ingesting files as of 0401 GMT on 4/26/01. Upon checking the MPL laptop there was a runtime error 75 on the screen and the MPL ingest program had been terminated. Closed out the error message @1425 GMT. Restarted the MPL ingest @1428 GMT. As of 1430 GMT the MPL appeared to be functioning properly and was ingesting files. RTS @ 1430 GMT.

4/19/01

CAM

SDS comlog reports indicated that the MPL had missed several collections. Check of instrument found the ingest process was hung up. Restarted the ingest process at 1800 GMT. The following collection was successful at 1801 GMT.

4/17/01

MK

Found the plastic cover still on top of the laser from the prior cleaning. Removed cover at 1430 GMT.

4/13/01

MR

The ingest has stopped. Rebooted the ingest at 1440 GMT. RTS at 1455 GMT.

4/5/01

JEM

There was a runtime 75 error displayed on the screen and the ingest has stopped. Restarted the ingest on the MPL. After restarting the ingest, the MPL scanned the directory and found 20 files and began to ingest them. After the files were done ingesting the MPL appeared to be functioning properly. RTS @1414 GMT.

4/2/01

MR

The SDS system was showing that the ingest was not working. T/S found the ingest was working. Shut down the ingest and the restarted from the desktop. Suspect that the computer had lost its link to the SDS. RTS at 1445 GMT.

3/29/01

MR

The ingest has stopped. Rebooted the ingest on the desktop. RTS at 1603.

3/26/01

MR

The ingest has stopped. Restarted the ingest at 1509 GMT. RTS at 1524.

3/16/01

MR

The ingest has stopped. Restarted the ingest from the desktop. RTS at 1532 GMT.

3/14/01

MR

The ingest has stopped. Restarted the ingest from the desktop. RTS at 1603.

3/12/01

KT

The laptop had lost it's link with the SDS. At 1841 GMT closed down the programs on the laptop and powered the laptop down. At 1843 GMT powered the laptop on and watched as the programs came up. At 1845 GMT after successful start up the ingest was operational. RTS at 1855 GMT.

3/8/01

MR

The ingest has stopped with a runtime error 75. Restarted the ingest at 1623 GMT. RTS at 1638.

3/7/01

MR

The ingest has stopped. Rebooted the computer at 1858 GMT. RTS at 1915 GMT.

3/5/01

MR

The ingest has stopped. Restarted the ingest. RTS at 1548 GMT.

2/28/01

MR

The collections have stopped. T/S found that the ingest was running but the link light on the network card was not on. Rebooted the computer and the light came back on. Collections were restored. RTS at 1708 GMT.

2/27/01

MR

The ingest has stopped. Rebooted the ingest at 1559. RTS at 1615 GMT.

2/26/01

MR

The laptop computer was found powered off. Power failure on 2/24/01 is suspected as causing the computer to shut down. Powered on the computer at 1527 GMT. System came up and started working at 1529. RTS at 1542 GMT.

2/21/01

MR

The program has terminated. Rebooted the computer at 1559. Program came up and started running at 1602 GMT. RTS at 1615 GMT.

2/20/01

MR

The ingest has stopped. Restarted the ingest at 1630 GMT. RTS 1645.

2/16/01

MR

The ingest has stopped. Restarted the ingest on the laptop. RTS at 1635 GMT.

2/13/01

MR

The ingest has stopped with a run time error 75. Shut the ingest down and then restarted the ingest with the software. RTS at 1525 GMT.

2/12/01

KT

The MPLHR ingest had stopped. At 1531 GMT closed ingest window and restarted ingest. At 1533 GMT ingest came on line and proceeded to ingest old records up to current. RTS at 1545 GMT.

2/8/01

MR

Rebooted the computer due to it not sending files to the SDS system. RTS at 1558 GMT.

2/7/01

MR

The SDS is not receiving any files from the MPLHR. Rebooted the computer at 1706 GMT. RTS at 1721 GMT.

2/6/01

MR

The SDS is not receiving files from the MPLHR. Rebooted the computer at 1535 GMT. RTS at 1550 GMT.

2/5/01

KT

The ingest had quit and needed to be restarted. At 1541 GMT closed the ingest window. At 1542 GMT the ingest was restarted. RTS at 1555 GMT.

1/30/01

MR

The ingest has stopped running. Restarted the ingest. RTS at 1530 GMT.

1/29/01

MR

The ingest has stopped. Restarted the ingest. RTS at 1525 GMT.

1/25/01

MR

The ingest has stopped. Rebooted the ingest with the software.

1/23/01

MR

The ingest had stopped with a "run time error 75" code. Restarted the ingest at 1636 GMT. RTS at 1650 GMT.

1/17/01

MR

The ingest has stopped. Restarted the ingest at 1555 GMT. RTS at 1610 GMT.

1/16/01

MR

The ingest has stopped. Restarted the ingest from the desktop. RTS at 2055 GMT.

1/16/01

MR

The ingest has stopped. Restarted the ingest. RTS at 1700 GMT.

1/10/01

MR

The ingest has stopped. Shut down the ingest and then restarted. RTS at 1510 GMT.

1/8/01

MR

The SDS is not able to collect data from the MPLHR. Observed that the ingest seemed to be working OK. Rebooted the computer at 1451 GMT. RTS at 1505. Collections was reestablished

1/4/01

MR

The ingest has stopped with a run time error 75. Restarted the ingest at 1523 GMT. RTS at 1538.

1/3/01

MR

The ingest has stopped. Restarted the ingest with the software. RTS at 1607 GMT.

12/29/00

KT

The was a file/path error 75 message. The ingest needed to be restarted. At 1627 GMT closed down the old ingest window. At 1630 GMT restarted ingest. At 1632 GMT ingest window came up and started ingesting where program left off. RTS at 1640 GMT.

12/21/00

MR

The collections has stopped with an error 75. Restarted the collections using the deskstop icon. RTS at 2130 GMT.

12/20/00

MR

The collections have stopped with a run time error of 75. Restarted the collections at 1532 GMT. RTS at 1547 GMT.

12/18/00

MR

The ingest has stopped. Restarted the ingest with the software at 1825 GMT. RTS at 1840 GMT.

12/11/00

KT

The ingest of data had stopped. At 1546 GMT observed Run time error '75', file/path access error. At 1548 GMT selected ok and the ingest window closed. At 1549 GMT restarted the ingest. At 1553 GMT ingest was running and previous data files were starting to get ingested. RTS at 1600 GMT.

12/8/00

MR

The ingest has stopped. Restarted the ingest. RTS at 1604.

12/07/00

KT

There was an MPL_ingest caution. Runtime error '75' Path/File access error. The ingest program had locked up. At 1544 GMT closed ingest window that was locked up and opened up new ingest window. At 1545 GMT ingesting of files began. RTS at 1557 GMT.

12/5/00

CAM

The instrument mentor sent a spare MPLHR *laser power supply to replace the failing original. Stopped the MPL program at 2050 GMT and powered the system down. Removed the old model 7300-L3 power supply and installed the new model R1-740 power supply. Powered the system on at 2015 GMT. Verified good readings and RTS status O.K. *Old - 3334 (WD27546) *New - 3689 (WD27547)

12/5/00

MR

The collections and ingest has failed with a "file/path access" error. Shut down the ingest and then restarted the ingest on the laptop. RTS at 1615.

12/4/00

MR

The collections had stopped with a run time error. Restarted the collections at 1545 GMT. RTS at 1600 GMT.

11/29/00

MR

The ingest has stopped. Restarted the ingest. RTS at 1616 GMT.

11/28/00

MR

The SDS has reported that the MPL has not sent any files to be ingested. The ingest on the MPL computer had stopped with a run time error. Restarted the ingest with the software. RTS at 1537.

11/27/00

MR

The SDS has reported that the MPL has not sent any files to be ingested. The ingest on the MPL computer had stopped with a run time error. Restarted the ingest with the software. RTS at 1245.

11/24/00

KT

During PM, the ingest displayed a run time error "75", Path/File access error. At 1522 GMT closed ingest window. At 1524 GMT opened ingest through shortcut. Ingest appeared and files began to ingest. RTS at 1535 GMT.

11/22/00

MR

The SDS has not collected any data from the MPL. Rebooted the MPLHR at 1534. RTS at 1549 GMT. Status O.K.

11/16/00

KT

The MPLHR laptop had no files to be ingested. At 2145 GMT observed that the laptop ingest had locked up. At 2146 GMT closed ingest window and restarted ingest through shortcut located on the desktop. At 2149 GMT the ingest started back up and downloaded last seven files that were not ingested. At 2153 GMT system was left with ingest working. RTS at 2200 GMT.

11/13/00

KT

There was no graphical data being displayed on the monitor. There was a runtime error 75 path/error message that was displayed upon shutdown of computer. There also was a problem with the collections, there were no files being sent to the SDS. At 1635 GMT brought the MPLHR laser power down to 0.0 watts and powered off. At 1637 GMT closed down the programs on the computer and powered the computer off. At 1640 GMT powered the whole system on. At 1644 GMT the system came on line. At 1701 GMT the SDS was able to collect files. RTS at 1705 GMT.

11/09/00

MR

The top graph was not being displayed. Rebooted the system at 1609 System came up and started running at 1612. RTS at 1624 GMT.

11/8/00

MR

The SDS has stopped receiving files from the SDS. Stopped the ingest process according to the procedure worked through with Conner Flynn. Restarted the ingest process with the software. System started to send files to the SDS. RTS at 1645

11/02/00

MR

The laser is not working. The emission light is out but the computer is still collecting data. Pressed the emission button and the light came on for about 2 seconds then went out. Pressed the button again and the emission light came on and stayed on. The laser also started working but the computer lost communications with the laser. Restarted the communications with the laser with the software on the computer. The computer is also not sending any data to the SDS system. Called Conner Flynn and Albert Mendosa and left a voice mail message to have them call me back so that they can walk me through a restart procedure for the data problem. RTS at 1525 GMT.

10/31/00

MR

The instrument stopped sending files to the SDS on 10/31/00 at 13:01 GMT. Shut the instrument down at 1518 GMT and checked all the data connections and the network connections. Reseated the PCMCIA cards and connections. Restarted the computer at 1524 GMT. System came up and started running at 1526 GMT. RTS at 1532.

10/30/00

MR

The SDS has not received any files from the MPL since 10/28/00 at 2201. Rebooted the computer. The system came up and started sending files to the SDS system. RTS at 1711 GMT.

10/27/00

JEM

The MPL has lost its link with the SDS again. Rebooted the MPL computer. When I checked the computer again after it came back online it was sending files. RTS @1421 GMT.

10/20/00

MR

The SDS has not ingested any data from the instrument since 10/19/00 at 4:01 GMT. Power cycled the instrument at 1422 GMT. System came up and started running at 1425 GMT and started to send files to the SDS. RTS at 1437 GMT.

10/13/00

MR

The MPLHR has not sent any files to the SDS. Rebooted the MPLHR computer. System started up at 1430 GMT. RTS at 1435 GMT.

10/10/00

MR

According to the SDS Ops page, the MPLHR has not collected since 10/5/00. Rebooted the computer. The computer started sending back files to the SDS that started on 10/5/00.

10/3/00

MR

The MPL missed the 2000 GMT collection. Power cycled the system at 2100 GMT. RTS at 2115.

5/24/00

MR

Power to the optical trailer is to be shut down for ½ the day. Gracefully shut down the computer at 1455. Powered back up at 2000.

5/23/00

MR

Instrument mentor replaced the *detector. It took a couple of more days to calibrate and setup system characterization. The system dark counts, after pulse with lid on and off, and overlap were all measured. The relevant data files have been provided to NASA GSFC. *New - 2165.

5/2/00

KT

At 1807 GMT mentor called and troubleshooting was performed to determine why no data was being displayed. At 1823 GMT mentor believed that the detector went bad. Waiting for mentor response before further troubleshooting takes place. RTS at 1824 GMT.

5/1/00

KT

At 1818 GMT observed that the cloud height graphs were not displaying any data. At 1821 GMT rebooted computer. There was no change. At 1831 GMT after talking with Chris Martin, shut down the instrument and computer. At 1833 GMT powered system back on. There was still no change. At 1855 GMT called and left message for mentor. Waiting for mentor reply. RTS at 1900 GMT.

4/18/00

MR

No data was being collected from the instrument. T/S indicated the AC power was out. Checked the UPS and found it dead. Power cycled the UPS and the AC power returned. System came up and started to take data at 1350.

4/7/00

MR

The computer was not sending files to the SDS. Rebooted at 1340. RTS at 1355.

4/3/00

KT

The computer needed to be rebooted due to network problems. At 1807 GMT rebooted computer. RTS at 1811 GMT.

3/7/2000

CAM

SDS comlog reports indicated that the MPLHR had not collected since 2204 GMT on 3/6. Instrument status check found the system was running with no obvious problems. Rebooted the laptop computer at 1415 GMT. Cleaned the hatch window and RTS status O.K. at 1500 GMT. SDS collections were successful at 1502 GMT.

3/3/2000

Mentor (Flynn)

Installed the repaired MPLHR.

1/28/00

GRS

The instrument Mentor was contacted on 1/24/00 about a problem with the instrument. It was not displaying any data. The Mentor thought that the laser diode power supply had gone bad and was in need of replacement. The laser diode power supply arrived and needed to be installed. R/R the laser diode power supply from 2125 to 2135 GMT. Powered up the instrument at 2137 and the computer at 2138 GMT. The program started at 2139. Data was only being plotted on one of the three graphs. That data was on the "clouds" graph and was a solid black line from top to bottom. Shut down the instrument and computer at 2202 GMT. Powered up the instrument and computer at 2206 GMT. The data started the same as the first time with the "clouds" graph being black from top to bottom. Called Albert Mendoza and Connor Flynn and left voice messages with both explaining the problem. Talked to Chris Martin and he informed me that is exactly what the instrument was doing before the laser diode power supply was replaced. So, the laser diode power supply does not seem to be the problem. Talked to Dave Breedlove and he had me shut down the system until we heard back from Connor or Albert.

1/24/00

KT

At 1901 GMT during rounds the computer was not displaying any graphical data. Called mentor and walked through and reseated the connections to the TTL output and Lidar power. Mentor called the detector bad. The system was powered down at 1955 GMT until     repairs can be made. RTS at 1955 GMT.

1/11/2000

KT

There was an error message stating that the controller was not responding. After rebooting attempts failed there was a loose cable connection. At 1955 GMT cable was seated tighter and system was rebooted. RTS at 1959 GMT.

1/6/2000

CM

While retiling the Optical trailer floor, contractors plugged a compressor into a UPS outlet which caused the breaker to trip. The UPS and all Optical trailer instruments lost power and had to be restarted. Removed the compressor from the UPS outlet and reset the breaker. Restarted all Optical trailer instruments. RTS status O.K.

01/03/00

KT

SDS reported that the MPLHR was not running. At 1931 GMT system was found operational. Rebooted system. RTS at 1937 GMT. Status ok.

12/28/99

CAM

The MPLHR laptop was not updating the data files. An error message was displayed indicating that the laser had stopped communicating or was powered off. Exited the program and rebooted the laptop at 1850 GMT. The data files then began updating normally. RTS status O.K. at 1900 GMT.

12/28/99

MK

Due to power line maintenance starting at 1500 GMT all instruments will need to be shut down. Shutdown MPLHR at 1435 GMT, restarted MPLHR at 1615 GMT. RTS at 1615 GMT.

11/1/99

KT

At 2058 GMT the MPLHR needed to be rebooted. RTS at 2107 GMT.

10/25/99

KT

At 1807 GMT there was an error message stating that the Controller was not responding or laser is off. Sampling process has stopped. At 1825 GMT the MPLHR was rebooted. RTS at 1825 GMT.

10/12/99

CAM

SDS comlog reports indicated that the MPLHR had not collected successfully since 2202 GMT on 10/10/99. The error message stated that an FTP problem occurred during the collection process. Instrument status check found an error message displayed on the MPLHR laptop. The message stated that the controller was not responding or the laser was turned off and that the sampling process was stopped. The last average time displayed on the monitor was 2248 GMT. UPS event logs also indicated that the site experienced a power bump at 2245 GMT on 10/10/99. OK'd the error message and rebooted the laptop at 1534 GMT. RTS status O.K. at 1535 GMT.

9/23/99

GRS

The onsite SDS operator indicated that we were not receiving files from the instrument. Troubleshooting found the GFCI receptacle that the instrument is plugged in to had tripped. The instrument shut down at 0549 GMT on the 23rd. Reset the GFCI receptacle at 1501 GMT. Shut down the program and computer at 1504 GMT. Power cycled the computer at 1505 GMT. The instrument started taking data at 1507 GMT. RTS

9/21/99

GRS

The MPLHR was found with part of it's program "hung up". The program clock was running but the data was not updating. The instrument went down on 9/20/99 at 0741 GMT. Shut the instrument down and rebooted the computer at 1535 GMT. The program came up normally and started taking data at 1537 GMT. The instrument had a good SDS collection at 1602 GMT. RTS

9/20/99

KT

At 2110 GMT checked status on MPLHR and error message stated that the laser was off. Clicked on ok and the laser data started scrolling again. RTS at 2120 GMT.

9/7/99

MR

The MPLHR computer was off, the laser was off and the laser power supply indicators where off. Checked the AC power plug and found the GFCI had been tripped. Reset the GFCI and the power was restored. Rebooted the computer and power cycled the laser power supply. The laser came back on and the computer rebooted successfully at 1449. Status O.K. RTS at 1500

8/2/99

MR

The laser controller had lost communications with the laptop. Checked the components and they all checked good. Restarted the data collections using the laptop software. Status O.K. RTS at 1515.

7/16/99

MR

The power had been reset to 1.0 watts last night by the equipment technician who was on site. The communications between the computer and the controller has been lost. Rebooted the MPLHR computer and the communications was restored to the controller and seemed to be working O.K.

7/1/99

MR

PM. The GFCI on the wall powering the laser power supply was tripped. Rest the circuit breaker and rebooted the computer. RTS at 1520.

6/30/99

MR

PM. The MPL has lost comms with the laser controller. Rebooted the MPLHR at 1405. Status O.K. The laser output is reading higher then set. It is reading .91 when set to .88. Reset the output to .86 in order to have the correct output of .88.

6/1/99

GRS

The laser had stopped lasing during the night. A large thunderstorm with a lot of lightning came over the site last night. Power cycled the MPL scaler and control box, and the laser diode module at 1633 GMT. Cleared the error message on the computer monitor. The MPL program did not come back completely. Rebooted the computer at 1653 GMT. The instrument came back on line at 1656 GMT. RTS

5/27/99

MR

PM. Removed the window and cleaned inside. RTS at 1410.

4/22/99

MR

Rebooted the MPLHR at 1456 due to power outage. RTS at 1510.

3/15/99

MW

The CF lost power due to a snow storm on 3/14/99 at 0030 GMT. The MPLHR has been down since that time. Powered the system back up after AC power was restored. RTS at 2034 GMT. Status OK.

2/23/99

CAM

SDS comlog report indicated that the MPLHR had not collected since 2258 GMT on 2/22/99. Status check found that the system was running and appeared normal. However, upgrades were made to the network cabling on 2/22/99 and the computer needs to be rebooted to reestablish the network connection. Rebooted the MPLHR computer at 1524 GMT on 2/23. MPLHR program restarted at 1527 GMT. RTS status O.K. at 1530 GMT. The following collection was successful at 1538 GMT.

1/25/99

MR

Instrument reinstalled by mentor. Status O.K.

12/7/98

SS

MPLHR has been removed from the site by mentor for repairs.

12/3/98

Mentor

Albert Mendoza is on site working on the instrument.

11/23/98

SS

Mentor requested that the instrument be shut down due to failure. Mentor is planning to visit the site for repairs. Instrument off line until further notice.

11/20/98

SS

The instrument mentor requested that the instrument be shut down. Adjusted laser power to 0.0 and turned off the MPL Scaler and Control, turned of the Laser Diode Module, shut down the laptop at 1550 GMT.

11/18/98

MW

The graphs on the laptop display are blank. The beam is transmitting and all indicators are OK. Called the instrument mentor and left a message. No corrective maintenance performed at this time.

11/17/98

ZW

The MPLHR did not display a graph. Rebooted the computer at 1630 GMT. The Graph is still not displayed at this time. RTS 1645 GMT.

11/12/98

SS

The laptop does not display a graph and the laptop laser power is 0.0 when the Spectra-Physics control panel indicates 1.01 watts. Called Conner Flynn and left a message explaining problem.

11/4/98

CM/TG

Instrument mentor supplied a replacement laptop for the MPLHR and requested that it be installed. Stopped the MPL program and powered off laptop at 1428 GMT. R/R the Panasonic *laptop with an IBM Thinkpad. Moved MPLHR from IP address 198.124.97.XX to 198.124.98.62 on hub #2/6. Powered the laptop on at 1448 GMT. Pinged the laptop and verified the network settings. RTS status O.K. at 1458 GMT. *Old - 7DKSA02304 *New - 97-65K2W

9/24/98

KT

There was a possible data interruption. The mentor request that the MPLHR laptop be re-booted. At 2124 GMT the laptop came back up. RTS system at 2125 GMT.

9/21/98

SS

The instrument mentor requested that the instrument be rebooted. Found the instrument not operating with an error on the laptop. Error read Tb2guest An error has occurred in your program. To keep working anyway , click Ignore and save your work in a new file. To quit this program, click Close. Clicked ignore and closed all programs and shut down computer then restarted it. RTS at 1422 GMT.

8/31/98

SS

Daily check revealed that the beam lens has insect droppings on lens. Cleaned lens with canned compressed gas. RTS at 1430 GMT.

8/26/98

SS

The laptop locked up 8/25/98 1552 GMT. Shut down the laptop and rebooted at 1511 GMT.

8/25/98

SS

Morning checks revealed that the computer was locked up and the date and time was 8/24/98 17:53. Used Control+Alt+Delete to bring up task manager and re-establish keyboard and mouse functions. reset time at 1505 GMT. System now operational. Rechecked the computer at 1530 GMT and it was locked up again. Repeated steps listed above to set the time at 1533 GMT.

8/24/98

SS

The time and date read 3 days and 6 hours off. Set the time at 1411 GMT.

8/20/98

MR

The time is 5 hr 22 min and 19 sec off. Reset the time. Status O.K.

8/19/98

MR

Connor Flynn called and asked if we could reboot the MPL. The laptop computer was locked up with a black screen and would not respond to any key strokes. Unplugged the power cord and the battery from the laptop. Plugged the power cord and the battery back into the computer and powered it up. The program started to run and seemed to be operating properly.

8/19/98

CAM

There was no data displayed on the MPL laptop computer. The display was blank as if the computer were turned off. Attempted to reboot using ctrl-alt-del with no success. The computer would not respond to any key commands. Attempts to power cycle the laptop also had no effect. The laptop was locked up tight. Removed all network and power connections from the laptop and removed the battery for 10 seconds. Replaced the battery and restored all connections. Powered the laptop on at 1600 GMT. During the boot up sequence a network error message was briefly displayed but the computer then booted up normally and began displaying data. No files have been received by SDS since 8/11/98. Mentor was on site at that time for maintenance and the computer was changed out with a laptop. Notified mentor of the situation. RTS at 1605 GMT.

8/19/98

MR

The time and date are wrong on the laptop computer display. Reset the time and date on the laptop using the DImision 4 program on the laptop. Time and date were showing 18:53 GMT 8/18/98. Time was corrected to 14:53 8/19/98.

8/18/98

MR

The display was not updating. Rebooted the system at 1445. System came up and started to run at 1458.

8/14/98

Mentor

Mentor installed MPL04. Instrument is high resolution version.

10/6/97

MW

The instrument mentor has removed the MPLHR from service and will send it to SESI for refocusing (under warranty). No corrective maintenance performed at this time.

9/15/97

MW

The MPLHR is down for unknown reasons. Mentor has the system off-line. No corrective maintenance performed.

9/3/97

CAM

Due to a Kay Electric transformer going down in Billings on 9/2, the site lost all power at around 2030 GMT. The extended power outage caused the optical trailer UPS to drain its batteries and all instruments lost power. When power was restored on 9/3 at 0142 GMT the MPLHR laser did not come up in the emission mode. Returned the Spectra-Physics keypad to emission mode at 1401 GMT on 9/3. System RTS status O.K. at 1402 GMT.

8/21/97

GRS

The Mentor could not talk to the instrument remotely through Timbuktu. T/S found that the Timbuktu program was in a "hung" state. Closed the Timbuktu program at 1610 GMT. Restarted the program at 1611 GMT. The instrument Mentor could then access the instrument. Instrument status O.K. RTS at 1615 GMT.

8/11/97

MW

The MPLHR computer was locked up. The display showed a blue screen which indicated a 'KERNEL_DATA_INPAGE_ERROR.' Rebooted computer at 1451 GMT. RTS at 1503 GMT.

7/28/97

MW

The MPLHR computer was not displaying its normal operational window. The screen was blue with several error messages displayed. Rebooted the computer at 1418 GMT. CHKDSK corrected a couple of disk errors. RTS at 1430 GMT.

7/28/97

CM/CW

AC power to the optical trailer had to be shut off at 1845 GMT for construction. Performed graceful shut down at 1830 GMT. AC power restored at 1905 GMT and system RTS status O.K. at 1905 GMT.

5/12/97

Mentor

Connor Flynn on site to install new MPL HR. This will replace the original failed unit that was taken off line 4/21/97.

4/30/97

ZW

Connor Flynn requested some measurements be made of the MPL HR Laser Telescope. Made the measurements and faxed them to Connor, also made a copy of measurements and gave the copy to Dan Nelson.

4/21/97

Mentor

Instrument experiencing problems and taken off line by mentor.

4/18/97

CAM

First system installed by Connor Flynn. System will be tested for deployment to TWP. Site Ops received operations training.